A bullet loss it's half of the velocity when its pass plank find the least number of plank

Grade:12th pass

1 Answers

Arun
5 years ago
Let   a = acceleration/deceleration of the bullet inside the plank.
let  s = thickness of the plank
let the initial velocity of the bullet before hitting the plank: = u
final velocity = v = u - u/n = u (n-1)/n
   v²  = u² + 2 a s
   u² (n-1)² / n² = u² + 2 a s
   2 a s = u² [ (n-1)² - n² ] / n²    =  - (2n - 1) u² / n²
   a = - (2 n - 1) u² / (n² 2 s)
let S be the thickness of the N planks kept one after another to stop the bullet.  The deceleration is the same.  The initial velocity is the same.
   v² = 0 = u² - 2 S * (2 n - 1) u² / (n² 2 s) 
           1 = S/s * (2 n  -1) / n²
  
    N=   S/s = n² / (2 n - 1)  =  1/2 * [ (2n² -n) + n ] / (2n -1)
               = 1/2 *  [ n + n / (2n -1) ]
               = n/2  +  n/(4n-2) 
               the second term above is between 0 and 1.
 N =  number of planks =    [ n/2 ] + 1
           where [ n/2 ]  is the greatest integer function for n/2.
 
Here n = 2
 
then
 
total number of planks = [2/2 ] + 1
 = 1 + 1 = 2